OpenStudy (anonymous):

am i doing this right?√-112 = 4√7

OpenStudy (asnaseer):

is that a minus sign within the square root?

OpenStudy (asnaseer):

i.e., is it \(\sqrt{-112}\)

OpenStudy (anonymous):

yes

OpenStudy (asnaseer):

then your answer needs to include \(i=\sqrt{-1}\)

OpenStudy (asnaseer):

i.e. it is a complex number

OpenStudy (anonymous):

yes and it has to be in i terms

OpenStudy (asnaseer):

e.g.: \(\sqrt{-8}=2\sqrt{2}i\) note that the \(i\) is outside the square root symbol

OpenStudy (anonymous):

then how do i do this one

OpenStudy (asnaseer):

you cn also write it as: \(2i\sqrt{2}\)

OpenStudy (asnaseer):

you have the right answer - you are just missing the 'i'.

OpenStudy (anonymous):

then do you just multipy it with the i to make it positive

OpenStudy (asnaseer):

"make it positive" ?

OpenStudy (anonymous):

well it is a negative number -112 so i put the i then √ to make it positive. I thought that ment i didnt need to use the i again

OpenStudy (asnaseer):

look at it this way:\[\sqrt{112}=4\sqrt{7}\]therefore:\[\sqrt{-112}\ne4\sqrt{7}\]otherwise you are saying that:\[\sqrt{112}=\sqrt{-112}\]

OpenStudy (asnaseer):

'i' is just a placeholder for \(\sqrt{-1}\)

OpenStudy (asnaseer):

so, for example:\[\sqrt{-4}=\sqrt{4}\times\sqrt{-1}=\sqrt{4}\times i=2i\]

OpenStudy (anonymous):

oic so would it be 4 √7i

OpenStudy (asnaseer):

yes - although I would probably write it as: \(4i\sqrt{7}\)

OpenStudy (asnaseer):

to make it clear that the i is not under the square root symbol

OpenStudy (anonymous):

oh so √-112 = 4i√7

OpenStudy (asnaseer):

thats is correct.

OpenStudy (anonymous):

thank you

OpenStudy (asnaseer):

yw :)